Grade 11Algebra

I) |4x-1|+|4x+5|=1
II) |x+1||x+4|=1
Please give me the answers with complete steps

Aniket
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Aniket
 
Always lerform these type type of question by breaking the interval. And also please ask only one question in one thread.
1. x
1 - 4x -4x -5 = 1
8x = -5
x = -5/8. (not accepeted)
Now
-5/4
1 -4x + 4x +5 = 1
6 = 1 (impossible)
Now
x> ¼
4x -1 + 4x +5 = 1
8x = -3
x = -3/8 (answer)
Hope it helps
 
Regards
Arun (askIITians forum expert)